2g-tm=5m-gh (solve for m)
Answer by checkley77(12844)

2g-tm=5m-gh (solve for m)
-tm-5m=-gh-2g
m(-t-5)=-g(h+2)
m=-g(h+2)/-(t+5)
m=g(h+2)/(t+5) ans.